A force of 5N is inclined at 45 degrees to a horizontal force of 8N. Find the magnitude of the resultant force and its direction?


1
Expert's answer
2021-09-01T11:20:18-0400

Sketch the diagram:



Find the x- and y-components:


"R_x=8+5\cos45=11.54\text N.\\
R_y=0+5\sin45=3.54\text N.\\
R=\sqrt{R_x^2+R_y^2}=12.1\text N.\\\space\\
\theta=\arctan\frac{R_y}{R_x}=17°."